# Explaining Tournament Solutions with Minimal Supports

### v2

@@ 3.1 MSs for Unweighted Tournaments
Abductive reasoning applied to tournament solutions explains the winner of a tournament by exhibiting a minimal subset of the tournament such that the winner remains unchanged. Let $G=(\mathcal{C},E)$ and $G^{\prime}=(\mathcal{C},E^{\prime})$ be two partial tournaments, we say that $G^{\prime}$ is an extension of $G$, denoted $G\subseteq G^{\prime}$, if $E\subseteq E^{\prime}$. If $E^{\prime}$ is complete, $G^{\prime}$ is called a completion of $G$, and we write $[G]$ for the set of completions of $G$. To obtain a formal definition we use the concept of necessary winner (Konczak and Lang 2005).
@@
Consider tournament $G$ in Figure 1(a). Both $\mathcal{X}$ and $\mathcal{Y}$ represented in Figure 2 are MSs for $a\in\operatorname{UC}(G)$. What $\mathcal{X}$ shows is that $a$ is not covered by $b$ since $a$ is preferred to $b$. $a$ is not covered by $c$ since $a$ is preferred to $b$ and $b$ is preferred to $c$. $a$ is not covered by $d$ since $a$ is preferred to $d$. Hence, $a$ is in the uncovered set. The case of $\mathcal{Y}$ is more compelling, as it immediately shows that $a$ is a Condorcet winner, i.e. is directly preferred to all its opponents.
@@ 3.2 MSs for Weighted Tournaments
Consider $G_{w}$ in Figure 1(b). As represented in Figure 3, both $\mathcal{I}$ and $\mathcal{J}$ are MSs for $a\in\operatorname{MM}(G_{w})$. While $\mathcal{I}$ shows that $a$ achieves a maximin score of 3 and that no other candidate can beat it since they all end up with three defeats against $a$, $\mathcal{J}$ only secures a maximin score of 2 and uses additional pairwise comparisons between candidates distinct from $a$ to show that the other candidates cannot outperform $a$.
@@ 3.3 Selecting among MSs for Explanations
For a given tournament there might be an exponential number of different minimal supports since MSs are an instance of a Sperner family (Sperner 1928). Therefore in this section we introduce criteria to select the most promising MSs from an explainability point of view.
Every MS can be a certificate for an explanation since by definition, it is subset minimal and present precisely enough information to guarantee the outcome of the tournament. Hence, looking back at the Gricean principles (Grice 1975) mentioned in the introduction, all MSs are optimal for quality. However, we can still use the manner criteria to discriminate between MSs since some can be smaller than others, i.e., rely on a smaller total amount of information. This is interesting from the perspective of formal verification since it is easier to verify the correctness of smaller certificates (imagine you have to call back participants or rewatch the games between contestants to verify that the registered preferences are correct).
To measure this amount of information in a general context, we take inspiration from the bribery setting and we fix the smallest atom of information to be the individual swap, i.e., the result of one pairwise comparison between two candidates for one participant.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$, and a winning candidate $w\in S(G)$. A MS $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in S(G)$ is a smallest minimal support (SMS) if and only if for all MS $\mathcal{Y}=(\mathcal{C},\mu_{\mathcal{Y}})$ for $w\in S(G)$, $\sum_{(c,c^{\prime})\in\mathcal{C}^{2}}\mu_{\mathcal{X}}(c,c^{\prime})\leq\sum_{(c,c^{\prime})\in\mathcal{C}^{2}}\mu_{\mathcal{Y}}(c,c^{\prime})$.
Smallest MSs for weighted tournaments can ve viewed as MSs such that no strictly smaller MS exists for the $\ell_{1}$-norm ($x||_{1}=\sum_{i\in I}|x_{i}|$) of the weight vector.
@@
Given a complete tournament $G=(\mathcal{C},\mu)$, a tournament solution $S$, and a winning candidate $w\in S(G)$, a shortest-paths SMS $\mathcal{X}$ for $w\in S(G)$, is a SMS such that for all $c\in\mathcal{C}\setminus{\{w\}}$, the directed path from $w$ to $c$ in $\mathcal{X}$ is the shortest in $G$.
For the myopic-score solutions, we argue that it is better to to put more emphasis on good performances of the winner rather than its opponents bad ones. Typically, if a candidate is a Condorcet winner, we would like to show it (see examples 1 and 2). Thus we will prioritise SMSs containing the highest number of won pairwise comparisons by the considered tournament winner.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$, a tournament solution $S$, and a winning candidate $w\in S(G)$. Given a SMS $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in S(G)$, we define the win count of $\mathcal{X}$ as $\operatorname{WC}(\mathcal{X})=\sum_{c\in\mathcal{C}}\mu_{X}(w,c)$. $\mathcal{X}$ is a maxwin-SMS if and only if for all SMS $\mathcal{Y}$ for $w\in S(G)$, $\operatorname{WC}(\mathcal{X})\geq\operatorname{WC}(\mathcal{Y})$.
@@ 4.1 Top Cycle and Uncovered Set
Top cycle ($\operatorname{TC}$) selects the smallest dominating set of candidate or in graph theory perspective, the top strongly-connected component. We use the alternative characterization based on paths to prove that all MSs for $\operatorname{TC}$ are directed trees.
@@
Given a complete tournament $G=(\mathcal{C},E)$, and a winning candidate $w\in\operatorname{TC}(G)$, for all MS $\mathcal{X}$ for $w\in\operatorname{TC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree, i.e., for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path from $w$ to $c$ in $\mathcal{X}$.
@@
Let $\mathcal{X}=(C,E_{\mathcal{X}})$ be a MS for $w\in\operatorname{TC}(G)$. Suppose there exists $c\in\mathcal{C}\setminus{\{w\}}$ such that there is no directed path from $w$ to $c$ in $\mathcal{X}$. Let $\mathcal{Y}$ be a completion of $\mathcal{X}$ such that $w$ and all the candidates reachable from $w$ in $\mathcal{X}$ lose their missing pair-wise comparisons. Then no new candidate is reachable from $w$ in $\mathcal{Y}$. Thus, there is no path in $\mathcal{Y}$ from $w$ to $c$. Hence, $w\not\in\operatorname{TC}(\mathcal{Y})$ and $w\not\in\text{NW}_{\operatorname{TC}}(\mathcal{X})$ and $\mathcal{X}$ is not a a SMS for $w\in\operatorname{TC}(G)$. Contradiction. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a directed path from $w$ to $c$ in $\mathcal{X}$.
Let us now show that for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path from $w$ to $c$ in $\mathcal{X}$. Suppose that exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that there are two directed paths $p^{\prime}$ and $p^{\prime\prime}$ from $w$ to $c^{*}$ in $\mathcal{X}$ such that $p^{\prime}$ ends with $(c^{\prime},c^{*})\in E_{\mathcal{X}}$, $p^{\prime\prime}$ ends with $(c^{\prime\prime},c^{*})\in E_{\mathcal{X}}$ and $c^{\prime}\neq c^{\prime\prime}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, we know there exists a directed path $p_{c}$ from $w$ to $c$ in $\mathcal{X}$. Let $\mathcal{Y}=(\mathcal{C},E_{\mathcal{Y}})$ be such that $E_{\mathcal{Y}}=E_{\mathcal{X}}\setminus{\{(c^{\prime\prime},c^{*})\}}$. If $p_{c}$ does not contains $(c^{\prime\prime},c^{*})$, $\mathcal{Y}$ also contains $p_{c}$. Else, if $p_{c}$ contains $(c^{\prime\prime},c^{*})$, $p_{c}$ can be decomposes as $p^{\prime\prime}+p_{c}^{\prime}$ where ++ is the concatenation and $p_{c}^{\prime}$ a path from $c^{*}$ to $c$ and $p^{\prime}+p_{c}^{2}$ is a directed path from $w$ to $c$ in $\mathcal{Y}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, exists a directed path from $w$ to $c$ in $\mathcal{Y}$, thus $w\in\text{NW}_{\operatorname{TC}}(\mathcal{Y})$ but $\mathcal{Y}\subsetneq\mathcal{X}$ so $\mathcal{X}$ is not a MS for $w\in\operatorname{TC}(G)$. Contradiction. ∎
From the common structure shared by MS for $\operatorname{TC}$, we can show that all MS and thus SMS have the same size.

### v3

@@ 3.1 MSs for Unweighted Tournaments
Abductive reasoning applied to tournament solutions explains the winner of a tournament by exhibiting a minimal subset of the tournament such that the winner remains unchanged. Let $G=(\mathcal{C},E)$ and $G^{\prime}=(\mathcal{C},E^{\prime})$ be two partial tournaments, we say that $G^{\prime}$ is an extension of $G$, denoted $G\subseteq G^{\prime}$, if $E\subseteq E^{\prime}$. If $E^{\prime}$ is complete, $G^{\prime}$ is called a completion of $G$, and we write $[G]$ for the set of completions of $G$. To obtain a formal definition we use the concept of necessary winner (konczak_voting_2005).
@@
Consider tournament $G$ in Figure 1(a). The partial tournament $\mathcal{X}$ represented in Figure 2(a) is an MS for $a\in\operatorname{UC}(G)$. Note that $\mathcal{X}\subseteq G$. What $\mathcal{X}$ shows is that $a$ is not covered by $b$ since $a$ is preferred to $b$; $a$ is not covered by $c$ since $a$ is preferred to $b$ and $b$ is preferred to $c$; $a$ is not covered by $d$ since $a$ is preferred to $d$. Hence, $a$ is in the uncovered set. Similar reasoning holds for $\mathcal{Y}$ in Figure 2(b).
@@ 3.2 MSs for Weighted Tournaments
Consider $G_{w}$ in Figure 1(b), both $\mathcal{I}$ and $\mathcal{J}$, represented in Figure 3, are MSs for $a\in\operatorname{MM}(G_{w})$. Observe that $\mathcal{I}\subseteq G_{w}$ and $\mathcal{J}\subseteq G_{w}$. $\mathcal{I}$ shows that $a$ achieves a maximin score of 3 and that no other candidate can beat it since they all end up with 3 defeats against $a$. $\mathcal{J}$ only secures a maximin score of 2 for $a$ and uses additional pairwise comparisons among other candidates to show that they cannot outperform $a$.
@@ 3.3 Selecting among MSs for Explanations
For a given tournament there might be an exponential number of different minimal supports since MSs are an instance of a Sperner family, a family of subsets which do not contain each others (sperner1928satz). Therefore in this section we introduce criteria to select the most promising MSs from an explainability point of view.
Every MS can be a certificate for an explanation since by definition it is subset minimal and presents precisely enough information to guarantee the outcome of the tournament. Hence, looking back at the Gricean principles (grice1975logic) mentioned in the introduction, all MSs are optimal for quality. However, we can still use the manner criteria to discriminate between MSs since some can be smaller than others, i.e., rely on a smaller total amount of information. This is interesting from the perspective of formal verification since it is easier to verify the correctness of smaller certificates (imagine you have to call back participants or re-watch the games between contestants to verify that the registered preferences are correct).
To measure this amount of information in a general context, we take inspiration from the bribery setting and we fix the smallest atom of information to be the individual swap, i.e., the preference of one voter in the pairwise comparison between two candidates.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$, and a winning candidate $w\in S(G)$, we define the size of an MS $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in S(G)$ as $|\mathcal{X}|=\sum_{(c,c^{\prime})\in\mathcal{C}^{2}}\mu_{\mathcal{X}}(c,c^{\prime})$. $\mathcal{X}$ is a smallest minimal support (SMS) for $w\in S(G)$ if and only if for all MSs $\mathcal{Y}$ for $w\in S(G)$, we have $|\mathcal{X}|\leq|\mathcal{Y}|$.
Example 4.
In Figure 2, $|\mathcal{X}|=3$, $|\mathcal{Y}|=3$ and both are SMSs for $a\in\operatorname{UC}(G)$. Both $\mathcal{I}$ and $\mathcal{J}$ in Figure 3 are MSs for $a\in\operatorname{MM}(G_{w})$. However, $\mathcal{I}$ is smaller than $\mathcal{J}$ as $|\mathcal{I}|=9$ and $|\mathcal{J}|=13$. Only $\mathcal{I}$ is an SMS for $a\in\operatorname{MM}(G_{w})$.
Smallest MSs for weighted tournaments can be viewed as MSs such that no strictly smaller MS exists for the $\ell_{1}$-norm ($x||_{1}=\sum_{i\in I}|x_{i}|$) of the tournament weight vector. For unweighted tournaments this reduces to minimizing the number of edges.
@@
Given a complete tournament $G=(\mathcal{C},\mu)$, a tournament solution $S$, and a winning candidate $w\in S(G)$, a shortest-path SMS $\mathcal{X}$ for $w\in S(G)$, is an SMS such that for all $c\in\mathcal{C}\setminus{\{w\}}$, the directed path from $w$ to $c$ in $\mathcal{X}$ is the shortest in $G$.
Example 5.
Looking back at Figure 2, the path from $a$ to $d$ is 1-long in $\mathcal{X}$ and 2-long in $\mathcal{Y}$ as it goes through $b$. Hence, we will prefer $\mathcal{X}$ with its shorter paths.
For the myopic-score solutions, we argue that it is better to put more emphasis on the good performances of the winner rather than its opponents bad ones. Typically, if a candidate is a Condorcet winner, we would like to show it like $\mathcal{I}$ in Figure 3(a). In this perspective, we will prioritize SMSs containing the highest number of won pairwise comparisons by the considered tournament winner.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$, a tournament solution $S$, and a winning candidate $w\in S(G)$. Given an SMS $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in S(G)$, we define the win count of $\mathcal{X}$ as $\operatorname{WC}(\mathcal{X})=\sum_{c\in\mathcal{C}}\mu_{X}(w,c)$. $\mathcal{X}$ is a maxwin-SMS if and only if for all SMSs $\mathcal{Y}$ for $w\in S(G)$, we have $\operatorname{WC}(\mathcal{X})\geq\operatorname{WC}(\mathcal{Y})$.
@@ 4 Computing SMS
In this section we present the technical results of this paper, characterizing the MSs and SMSs for common tournament solutions, and providing polynomial algorithms for their computation when this is possible. All missing proofs can be found in the Appendix of this paper.
@@ 4.1 Top Cycle and Uncovered Set
Top cycle ($\operatorname{TC}$) selects the smallest dominating set of candidates, or, in graph-theoretic perspective, the top strongly-connected component. Here we use an alternative characterization based on paths to prove that all MSs for $\operatorname{TC}$ are directed trees.
@@
Given a complete tournament $G=(\mathcal{C},E)$, and a winning candidate $w\in\operatorname{TC}(G)$, for all MSs $\mathcal{X}$ for $w\in\operatorname{TC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree, i.e., for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path from $w$ to $c$ in $\mathcal{X}$.
@@
Let $\mathcal{X}=(C,E_{\mathcal{X}})$ be an MS for $w\in\operatorname{TC}(G)$. Suppose there exists $c\in\mathcal{C}\setminus{\{w\}}$ such that there is no directed path from $w$ to $c$ in $\mathcal{X}$. Let $\mathcal{Y}$ be a completion of $\mathcal{X}$ such that $w$ and all the candidates reachable from $w$ in $\mathcal{X}$ lose their missing pairwise comparisons, i.e. such that $\forall(u,v)\in R_{w}\times\mathcal{C}$, $(u,v)\in\mathcal{Y}$ implies $(u,v)\in\mathcal{X}$ or $v\in R_{w}$ where $R_{w}$ is the set of reachable candidates from $w$. Then no new candidate is reachable from $w$ in $\mathcal{Y}$. Thus, there is no path in $\mathcal{Y}$ from $w$ to $c$. Hence, $w\not\in\operatorname{TC}(\mathcal{Y})$ and $w\not\in\text{NW}_{\operatorname{TC}}(\mathcal{X})$ and $\mathcal{X}$ is not a an SMS for $w\in\operatorname{TC}(G)$. Contradiction. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a directed path from $w$ to $c$ in $\mathcal{X}$.
Let us now show that for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path from $w$ to $c$ in $\mathcal{X}$. Suppose that there exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that there are two directed paths $p^{\prime}$ and $p^{\prime\prime}$ from $w$ to $c^{*}$ in $\mathcal{X}$ such that $p^{\prime}$ ends with $(c^{\prime},c^{*})\in E_{\mathcal{X}}$, $p^{\prime\prime}$ ends with $(c^{\prime\prime},c^{*})\in E_{\mathcal{X}}$ and $c^{\prime}\neq c^{\prime\prime}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, we know there exists a directed path $p_{c}$ from $w$ to $c$ in $\mathcal{X}$. Let $\mathcal{Y}=(\mathcal{C},E_{\mathcal{Y}})$ be such that $E_{\mathcal{Y}}=E_{\mathcal{X}}\setminus{\{(c^{\prime\prime},c^{*})\}}$. If $p_{c}$ does not contain $(c^{\prime\prime},c^{*})$, $\mathcal{Y}$ also contains $p_{c}$. Else, if $p_{c}$ contains $(c^{\prime\prime},c^{*})$, $p_{c}$ can be decomposed as $p^{\prime\prime}\circ p_{c}^{\prime}$ where $\circ$ is the concatenation and $p_{c}^{\prime}$ a path from $c^{*}$ to $c$. $p^{\prime}\circ p_{c}^{\prime}$ is a directed path from $w$ to $c$ in $\mathcal{Y}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a directed path from $w$ to $c$ in $\mathcal{Y}$, thus $w\in\text{NW}_{\operatorname{TC}}(\mathcal{Y})$ but $\mathcal{Y}\subsetneq\mathcal{X}$ so $\mathcal{X}$ is not an MS for $w\in\operatorname{TC}(G)$. Contradiction. ∎
From the common structure shared by MSs for $\operatorname{TC}$, we can show that all MSs and thus SMSs have the same size.
Theorem 2.
Finally, computing a shortest-path SMS for $\operatorname{TC}$ can be achieved in polynomial time with Dijkstra’s algorithm (dljkstra1959note).
@@
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{TC}(G)$, there exists an algorithm which computes a shortest-path SMS for $w\in\operatorname{TC}(G)$ in $\mathcal{O}(m^{2})$.
Like $\operatorname{TC}$, the uncovered set ($\operatorname{UC}$) has a characterization based on the existence of directed paths between the winner and the other candidates, except that these paths have a maximum length of two. Hence, the following results can be derived from the previous propositions.
Corollary 4.
Given a complete tournament $G=(\mathcal{C},E)$, and a winning candidate $w\in\operatorname{UC}(G)$, for all MSs $\mathcal{X}$ for $w\in\operatorname{UC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree of depth at most 2, i.e., for alls $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path of length at most 2 from $w$ to $c$ in $\mathcal{X}$.
Corollary 5.
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{UC}(G)$, for all MSs $\mathcal{X}$ for $w\in\operatorname{UC}(G)$, we have $|\mathcal{X}|{=}m-1$.
Corollary 6.
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{UC}(G)$, there exists an algorithm which computes a shortest-path SMS for $w\in\operatorname{UC}(G)$ in $\mathcal{O}(m^{2})$.
4.2 Weighted Uncovered Set
To construct SMSs for the weighted uncovered set ($\operatorname{wUC}$), one can generalize the alternative characterization of the uncovered set to the case of weighted tournaments given by doring2023margin.
Lemma 7 (Lemma 2.2 in doring2023margin doring2023margin).
From Lemma 7, we obtain a result similar to Corollary 4 on the structure of any minimal support for $\operatorname{wUC}$.
Proposition 8.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, a winning candidate $w\in\operatorname{wUC}(G)$, for all MSs $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in\operatorname{wUC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree of depth at most 2 such that for all losing candidates $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil$ or there exists a unique $c^{\prime}$ such that $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$.
Even if MSs for $\operatorname{wUC}$ follow a structure similar to the unweighted case, we now show that deciding whether there exists a MS smaller than a given size is NP-complete.
Theorem 9.
Given $n\geq 2$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, a winning candidate $w\in\operatorname{wUC}(G)$, and an integer $k$, determining if there exists an MS of size at most $k$ is NP-complete.
Proof sketch.
To prove NP membership, note that Proposition 8 gives an easy to verify structure to identify if a partial tournament is an MS for $\operatorname{wUC}$.
To prove NP-hardness, we proceed with a reduction from the NP-complete Set Cover problem (karp1972reducibility). Given an instance of set cover problem with a universe $U$ of $p$ elements and a set $S$ of $q$ subsets of $U$, we show that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists an MS for $\operatorname{wUC}$ of size less than $k^{\prime}=p+q+k$ for a designated winner in a specific tournament with candidates associated to each element of $U$ and $S$. ∎
Even if computing SMSs for $\operatorname{wUC}$ is a hard problem, we provide tight bounds on their size.
Proposition 10.
Given $n\geq 2$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m\geq 3$, and a winning candidate $w\in\operatorname{wUC}(G)$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$, we have $n+m-2\leq|\mathcal{X}|\leq(n+1)(m-1)$.
4.3 Maximin
The maximin rule ($\operatorname{MM}$) selects the candidates with the least worst performance in a head-to-head against another alternative. Thus, an MS for $\operatorname{MM}$ must guarantee overall good performances for the winner while ensuring that the other candidates perform poorly at least once.
Lemma 11.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, and a winning candidate $w\in\operatorname{MM}(G)$ with a maximin score of $\sigma_{w}$, for all MSs $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in\operatorname{MM}(G)$, there exists $t\leq\left\lceil n/2\right\rceil$ such that, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq t$ and there exists $c^{\prime}\in\mathcal{C}$ such that $\mu_{\mathcal{X}}(c^{\prime},c)\geq n-t$.
From the previous characterization, we can derive a closed-form expression for the size of an SMS for $\operatorname{MM}$.
Theorem 12.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{MM}(G)$ with a maximin score of $\sigma_{w}$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{MM}(G)$, we have $|\mathcal{X}|{=}n(m-1)-t|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$ where $t=\min(\sigma_{w},\left\lfloor n/2\right\rfloor)$.